On January 1, 2021, Byner Company purchased a used tractor. Byner paid $4,000 down and signed a noninterest-bearing note requiring $33,000 to be paid on December 31, 2023. The fair value of the tractor is not determinable. An interest rate of 12% properly reflects the time value of money for this type of loan agreement. The company’s fiscal year-end is December 31. (FV of $1PV of $1FVA of $1PVA of $1FVAD of $1 and PVAD of $1(Use appropriate factor(s) from the tables provided.)
 
Required:
1. Prepare the journal entry to record the acquisition of the tractor.
2. How much interest expense will the company include in its 2021 and 2022 income statements for this note?
3. What is the amount of the liability the company will report in its 2021 and 2022 balance sheets for this note?
Answer #1

Req 1) use PV of $1

                                      (12%,3yrs)

Step 1) 4,000+(33,000*0.71178)= 27,488.74

dr. tractor                                   27,489

dr discount on notes payable     9,511

   cr. cash                                                4,000

   cr. notes payable                                  33,000

step 2) 33,000+4,000-27,489=9,511


Req 2 and 3)

interest expense 2021: (33,000*0.71178)*12%= 2,818.6488 or 2,819

(33,000*0.71178)= 23,488.74

interest expense 2022: (23,489+2,819)*12%= 3,156.96 or 3,157


liability amount 2021: 23,489+2,819= 26,308

liability amount 2022: 26,308+3,157= 29,465


Answer:


20212022
interest expense2,8193,157
liability amount26,30829,465


answered by: anonymous